# FinTagging: An LLM-ready Benchmark for Extracting and Structuring Financial Information

### v1

@@ 3.4.3 Evaluation setting
We use the LM Evaluation Harness eval-harness to build customized benchmark suites. Proprietary models such as GPT are accessed via the OpenAI API, while all other models, including DeepSeek and open-source LLMs, are evaluated through the TogetherAI API without local deployment. We standardize the input length to 2,048 tokens for FinNI and 4,096 for FinCL, with a generation limit of 1,024 tokens to support reasoning-intensive outputs. For fine-tuning strong PTMs such as BERT-large, FinBERT, and SECBERT, we curate 10 additional annual reports as training data (see Appendix F). All data follow the FiNER loukas2022finer and FNXL sharma2023financial formats. For retrieval in FinCL, we use ElasticSearch [^8] https://www.elastic.co/ with text-embedding-3-small as the embedding model.
4 Experiment and Result
4.1 Overall Results
@@
Table 5 presents the overall performance on the FinTagging benchmark. It clearly demonstrates that under our framework, LLMs can effectively handle both frequent and rare financial tags, indicating their ability to mitigate long-tail label challenges and underscoring the advantage of our information extraction and alignment formulation over traditional token-level classification approaches.
@@
Table 5: Overall Performance. Bolded values indicate the best performance, underlined values represent the second-best, and italicized values denote the third-best performance.
@@
Closed-source LLM GPT-4o 0.0764 0.0576 0.0508 0.0947 0.0788 0.0860
Open-source LLMs DeepSeek-V3 0.0813 0.0696 0.0582 0.1058 0.1217 0.1132
DeepSeek-R1-Distill-Qwen-32B 0.0482 0.0288 0.0266 0.0692 0.0223 0.0337
Qwen2.5-14B-Instruct 0.0423 0.0256 0.0235 0.0197 0.0133 0.0159
gemma-2-27b-it 0.0430 0.0273 0.0254 0.0519 0.0453 0.0483
Llama-3.1-8B-Instruct 0.0287 0.0152 0.0137 0.0462 0.0154 0.0231
Llama-3.2-3B-Instruct 0.0182 0.0109 0.0083 0.0151 0.0102 0.0121
Qwen2.5-1.5B-Instruct 0.0180 0.0079 0.0069 0.0248 0.0060 0.0096
Qwen2.5-0.5B-Instruct 0.0014 0.0003 0.0004 0.0047 0.0001 0.0002
Financial LLM Fino1-8B 0.0299 0.0146 0.0140 0.0355 0.0133 0.0193
Fine-tuned PLMs BERT-large 0.0135 0.0200 0.0126 0.1397 0.1145 0.1259
FinBERT 0.0088 0.0143 0.0087 0.1293 0.0963 0.1104
SECBERT 0.0308 0.0483 0.0331 0.2144 0.2146 0.2145
From the macro-level perspective, which emphasizes balanced performance across both frequent and rare tags, DeepSeek-V3 and GPT-4o achieve the highest macro-F1 scores (0.0582 and 0.0508), outperforming all fine-tuned PLMs. This highlights the strong generalization of large LLMs and the effectiveness of our task design. DeepSeek-R1-Distill-Qwen-32B also achieves a solid macro-F1 (0.0266), suggesting that good architecture and pretraining can help smaller models perform well in zero-shot settings. From the micro-level perspective, which favors frequent labels, DeepSeek-V3 again performs strongly with a micro-F1 of 0.1132, ranking third overall despite no fine-tuning. GPT-4o also performs competitively with a score of 0.0860, outperforming most open-source and domain-specific models.
4.2 Subtask Results
4.2.1 FinNI subtask
Table 6 shows the performance of various LLMs on the FinNI subtask. Overall, larger models perform better at identifying numerical entities and generating structured outputs, even without financial domain training. DeepSeek-V3 achieves the highest precision and recall, outperforming all other models. In contrast, smaller models like Qwen2.5-1.5B and -0.5B score below 0.1 F1, highlighting their limitations in zero-shot settings. Interestingly, DeepSeek-R1-Distill-Qwen-32B shows strong precision, suggesting that model design and distillation can help mitigate the impact of smaller scale.
Table 6: Performance comparison of different models on the FinNI subtask. Bolded values indicate the best performance, underlined values represent the second-best, and italicized values denote the third-best performance.
@@ 4.2.1 FinNI subtask
Closed-source LLM GPT-4o 0.6105 0.5941 0.6022
Open-source LLMs Deepseek-V3 0.6329 0.8452 0.7238
DeepSeek-R1-Distill-Qwen-32B 0.5490 0.2238 0.3180
Qwen2.5-14B-Instruct 0.3632 0.0018 0.0035
gemma-2-27b-it 0.5319 0.5490 0.5403
Llama-3.1-8B-Instruct 0.3346 0.1746 0.2295
Llama-3.2-3B-Instruct 0.1887 0.1794 0.1839
Qwen2.5-1.5B-Instruct 0.1323 0.0636 0.0859
Qwen2.5-0.5B-Instruct 0.0116 0.0027 0.0043
Financial LLM Fino1-8B 0.3416 0.1481 0.2066
However, the domain-specific Fino1-8B model, fine-tuned on the financial reasoning QA task, does not perform competitively. This indicates that only task-specific training can effectively improve performance on FinNI; pretraining on financial corpora alone offers limited benefit if the task is not well aligned.
4.2.2 FinCL subtask
Table 7: Performance comparison of different models on the FinCL subtask. Bolded values indicate the best performance, underlined values represent the second-best, and italicized values denote the third-best performance.
@@ 4.2.2 FinCL subtask
Closed-source LLM GPT-4o 0.1664
Open-source LLMs Deepseek-V3 0.1715
DeepSeek-R1-Distill-Qwen-32B 0.1013
Qwen2.5-14B-Instruct 0.1072
gemma-2-27b-it 0.1009
Llama-3.1-8B-Instruct 0.0807
Llama-3.2-3B-Instruct 0.0375
Qwen2.5-1.5B-Instruct 0.0419
Qwen2.5-0.5B-Instruct 0.0246
Financial LLM Fino1-8B 0.0704
Table 7 shows the accuracy of different models on the FinCL subtask. Overall performance is low, highlighting the challenge of fine-grained concept linking in finance. DeepSeek-V3 achieves the highest accuracy (0.1715), followed by GPT-4o, while all other models score below 0.11. Even large open-source models like Qwen2.5-14B and Gemma-2-27B underperform, and smaller models perform near random. This reflects the difficulty of handling complex taxonomies and subtle financial semantics.
4.3 Ablation analysis
[…]
Table 8: Performance comparison between w/wo our evaluation framework on the FinTagging benchmark dataset.
@@ 4.3 Ablation analysis
FinTagging GPT-4o 0.1075 0.0972 0.1021
Deepseek-V3 0.1229 0.1519 0.1359
Fino1-8B 0.0283 0.0138 0.0185
@@
Table 8 shows that all LLMs fail completely with extreme classification, yielding zero precision, recall, and F1, confirming that a single-step choice among thousands of flat labels is not a valid LLM evaluation protocol. By decoupling extraction from concept linking and covering the full 10k+ taxonomy, FinTagging produces meaningful scores and thus offers a far more realistic test bed for future model improvements.
5 Conclusion
@@
This paper presents FinTagging, a benchmark for evaluating large language models on XBRL tagging of real-world financial reports. The task is divided into two subtasks, financial numeric identification (FinNI) and concept linking (FinCL), to enable fine-grained evaluation of both information extraction and taxonomy alignment. Results show that while LLMs generalize well to long-tail entities and perform competitively in zero-shot settings, they struggle with accurate alignment to GAAP concepts. This reveals limitations in schema-aware reasoning and highlights the need for better semantic understanding. FinTagging offers a foundation for advancing research in XBRL tagging and regulatory reporting. The limitations of our work and directions for future research are discussed in Appendix A.

### v2

@@ 3.6.3. Evaluation setting
We use the LM Evaluation Harness (Gao et al., 2024) to build customized benchmark suites. Proprietary models are accessed via APIs, while open-source models are evaluated locally on a 4×H100 GPU cluster (80GB each). We standardize the input length to 2,048 tokens for FinNI and 4,096 for FinCL, with a generation limit of 1,024 tokens to support reasoning-intensive outputs. For fine-tuning strong PTMs such as BERT-large, FinBERT, and SECBERT, we curate 10 additional annual reports as training data (see Appendix I). All data follow the FiNER (Loukas et al., 2022) and FNXL (Sharma et al., 2023) formats. For retrieval in FinCL, we use ElasticSearch [^8] https://www.elastic.co/ with text-embedding-3-small as the embedding model. For evaluation, we spent a total of 500 GPU hours and about \$1,500 for GPT (from OpenAI) and DeepSeek (from TogetherAI) API.
4. Experiment and Result
4.1. Overall Results
@@
Table 5 presents the overall performance on the 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863i0.43922 0.45098 0.95294n0.5098 0.43529 0.93725T0.58431 0.42353 0.92157a0.6549 0.40784 0.90588g0.72549 0.39216 0.8902g0.79608 0.37647 0.87451i0.87059 0.36471 0.85882n0.94118 0.34902 0.84314g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: benchmark. It clearly demonstrates that under our framework, LLMs can effectively handle both frequent and rare financial tags, indicating their ability to mitigate long-tail label challenges and underscoring the advantage of our information extraction and alignment formulation over traditional token-level classification approaches.
@@
Table 5. Overall Performance. Bolded values indicate the best performance, underlined values represent the second-best, and italicized values denote the third-best performance.
@@
Closed-source LLM GPT-4o 0.0865 0.0626 0.0569 0.1052 0.0731 0.0863
Open-source LLMs DeepSeek-V3 0.0949 0.0778 0.0659 0.1074 0.1264 0.1162
Llama-4-Scout-17B-16E-Instruct 0.0683 0.0414 0.0400 0.1045 0.0526 0.0699
Llama-3.3-70B-Instruct 0.0544 0.0279 0.0288 0.0665 0.0382 0.0485
DeepSeek-R1-Distill-Qwen-32B 0.0532 0.0283 0.0285 0.0814 0.0214 0.0339
Qwen3-32B 0.0639 0.0314 0.0324 0.1127 0.0230 0.0382
gemma-2-27b-it 0.0471 0.0291 0.0276 0.0533 0.0390 0.0451
Qwen3-14B 0.0591 0.0274 0.0288 0.1069 0.0182 0.0311
Llama-3.1-8B-Instruct 0.0345 0.0178 0.0169 0.0575 0.0166 0.0258
Llama-3.2-3B-Instruct 0.0194 0.0118 0.0100 0.0178 0.0084 0.0114
Qwen3-1.7B 0.0207 0.0067 0.0080 0.1063 0.0031 0.0060
Qwen3-0.6B 0.0026 0.0008 0.0010 0.0562 0.0002 0.0005
Financial LLM Fino1-8B 0.0344 0.0143 0.0151 0.0419 0.0128 0.0197
Fine-tuned PLMs BERT-large 0.0252 0.0266 0.0205 0.1518 0.1283 0.1391
FinBERT 0.0046 0.0064 0.0042 0.0872 0.0526 0.0656
SECBERT 0.0231 0.0295 0.0203 0.1870 0.1697 0.1779
From the macro-level perspective, which emphasizes balanced performance across both frequent and rare tags, DeepSeek-V3, GPT-4o, and Llama-4-Scout-17B-16E-Instruct achieve the top 3 macro-F1 scores (0.0659, 0.0569, and 0.0400), outperforming all fine-tuned PLMs. This highlights the strong generalization of large LLMs and the effectiveness of our task design. Qwen3-32B also achieves a solid macro-F1 (0.0324), suggesting that good architecture and pretraining can help smaller models perform well in zero-shot settings. From the micro-level perspective, which favors frequent labels, DeepSeek-V3 again performs strongly with a micro-F1 of 0.1162, ranking third overall despite no fine-tuning. GPT-4o also performs competitively with a score of 0.0863, outperforming most open-source and domain-specific models.
4.2. Subtask Results
4.2.1. FinNI subtask
Table 6 presents the results of different models on the FinNI subtask. Overall, larger models demonstrate stronger capabilities in identifying numerical entities and producing structured outputs, even without financial domain adaptation. Among them, DeepSeek-V3 achieves the best performance with both the highest recall (0.8430) and F1 (0.6932), while also maintaining competitive precision. In comparison, GPT-4o, the closed-source baseline, performs relatively well with balanced scores, securing the second-best F1 (0.5397). Interestingly, Qwen3-32B and Qwen3-14B yield very high precision (0.6991 and 0.6912, respectively), but their recall is extremely low, leading to weak F1 scores, which highlights that precision alone is insufficient in this task.
Table 6. Performance comparison of different models on the FinNI subtask. Bolded values indicate the best performance, underlined values represent the second-best, and italicized values denote the third-best performance.
@@ 4.2.1. FinNI subtask
Closed-source LLM GPT-4o 0.5893 0.4977 0.5397
Open-source LLMs Deepseek-V3 0.5886 0.8430 0.6932
Llama-4-Scout-17B-16E-Instruct 0.4668 0.3164 0.3771
Llama-3.3-70B-Instruct 0.4826 0.3301 0.3920
DeepSeek-R1-Distill-Qwen-32B 0.5676 0.1942 0.2894
Qwen3-32B 0.6991 0.1804 0.2868
gemma-2-27b-it 0.5060 0.4526 0.4778
Qwen3-14B 0.6912 0.1487 0.2448
Llama-3.1-8B-Instruct 0.3874 0.1761 0.2421
Llama-3.2-3B-Instruct 0.1856 0.1203 0.1460
Qwen3-1.7B 0.7362 0.0281 0.0541
Qwen3-0.6B 0.2803 0.0019 0.0038
Financial LLM Fino1-8B 0.3431 0.1293 0.1878
Smaller open-source models such as Qwen3-1.7B and Qwen3-0.6B show severe limitations, with F1 below 0.1, underscoring the difficulty of the FinNI subtask in zero-shot settings. In contrast, models like gemma-2-27B-it achieve relatively balanced results, ranking third in both recall and F1. Finally, the domain-specific Fino1-8B, despite being fine-tuned on financial reasoning QA, underperforms compared to general-purpose LLMs. This suggests that domain pretraining on financial text alone provides limited benefit unless the training task aligns closely with FinNI’s requirements.
4.2.2. FinCL subtask
Table 7. Performance comparison of different models on the FinCL subtask. Bolded values indicate the best performance, underlined values represent the second-best, and italicized values denote the third-best performance.
@@ 4.2.2. FinCL subtask
Closed-source LLM GPT-4o 0.1829
Open-source LLMs Deepseek-V3 0.1889
Llama-4-Scout-17B-16E-Instruct 0.1649
Llama-3.3-70B-Instruct 0.1318
DeepSeek-R1-Distill-Qwen-32B 0.1141
Qwen3-32B 0.1277
gemma-2-27b-it 0.1099
Qwen3-14B 0.1144
Llama-3.1-8B-Instruct 0.0913
Llama-3.2-3B-Instruct 0.0415
Qwen3-1.7B 0.0735
Qwen3-0.6B 0.0414
Financial LLM Fino1-8B 0.0807
Table 7 reports the accuracy of different models on the FinCL subtask, which is performed on candidate sets obtained from the best top-200 retrieval results, as shown in Table 16 (Appendix J). Although the retrieval stage provides a foundation, the low Acc@k values, especially for table-based entities, directly constrain the FinCL performance.
Within this reranking formulation, DeepSeek-V3 achieves the best accuracy at 0.1889, closely followed by GPT-4o at 0.1829 and Llama-4-Scout-17B-16E-Instruct at 0.1649. Other large open-source models perform notably worse, with Llama-3.3-70B at 0.1318. The medium-scale models, such as Qwen3-32B (0.1277), DeepSeek-R1-Distill-Qwen-32B (0.1141), gemma-2-27B-it (0.1099), and Qwen3-14B (0.1144), show similar levels. While smaller models, including Llama-3.2-3B and Qwen3-0.6B, perform near random with 0.0415 and 0.0414. The domain-specific Fino1-8B also lags at 0.0807, suggesting that domain pretraining without explicit task alignment offers little benefit. As a result, overall accuracy remains low, reflecting both the inherent difficulty of handling complex taxonomies and subtle financial semantics, and the cascading effect of retrieval errors that further lowers FinCL performance.
4.3. Ablation analysis
We further compare our benchmark against extreme multi-class classification settings. The prompt template is shown in Figure 11 in the Appendix, and we select the best-performing model from each category (closed-source, open-source, and financial LLMs). The evaluation uses triplet-level ($\{\texttt{Tag},\texttt{Fact},\texttt{Type}\}$) Precision, Recall, and F1.
Table 8. Performance comparison between with/without our evaluation framework on the 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863i0.43922 0.45098 0.95294n0.5098 0.43529 0.93725T0.58431 0.42353 0.92157a0.6549 0.40784 0.90588g0.72549 0.39216 0.8902g0.79608 0.37647 0.87451i0.87059 0.36471 0.85882n0.94118 0.34902 0.84314g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: benchmark dataset.
@@ 4.3. Ablation analysis
FinTagging GPT-4o 0.1166 0.0915 0.1026
Deepseek-V3 0.1187 0.1581 0.1356
Fino1-8B 0.0336 0.0140 0.0197
@@
Table 8 shows that extreme classification is not a valid way to evaluate LLMs on XBRL tagging. This single-step formulation requires presenting the full space of thousands of US-GAAP concepts and directly assigning tags to raw text, without first identifying which spans are true numerical entities. In practice, this is infeasible: the model cannot realistically process the entire label space and is forced to rely only on its internal knowledge, which leads to meaningless zero-shot predictions. As the results indicate, all tested models, including GPT-4o and DeepSeek-V3, collapse to zero precision, recall, and F1 under this setting.
In contrast, our 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863i0.43922 0.45098 0.95294n0.5098 0.43529 0.93725T0.58431 0.42353 0.92157a0.6549 0.40784 0.90588g0.72549 0.39216 0.8902g0.79608 0.37647 0.87451i0.87059 0.36471 0.85882n0.94118 0.34902 0.84314g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: benchmark reformulates the task into a two-stage process that aligns with how tagging is actually performed. The FinNI first detects and types coarse-grained numerical entities from text and tables, ensuring that only valid values enter the linking step. The FinCL then narrows the taxonomy search space by retrieving a candidate set of relevant tags and requiring the LLM to select the best match using contextual reasoning. With this design, models achieve non-trivial performance: for example, DeepSeek-V3 reaches an F1 of 0.1356 and GPT-4o 0.1026. These results demonstrate that FinTagging provides a more faithful evaluation of LLMs’ zero-shot tagging capabilities, surfacing meaningful signals where extreme classification fails.
4.4. Error propagation in 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863i0.43922 0.45098 0.95294n0.5098 0.43529 0.93725T0.58431 0.42353 0.92157a0.6549 0.40784 0.90588g0.72549 0.39216 0.8902g0.79608 0.37647 0.87451i0.87059 0.36471 0.85882n0.94118 0.34902 0.84314g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: pipeline
Table 9 breaks down model accuracy across three evaluation settings: full pipeline, FinCL, and reranking-only in FinCL with guaranteed gold concepts in the top 200 candidates. The results reveal clear evidence of cascading errors. End-to-end pipeline accuracy is consistently the lowest, since errors from entity extraction (FinNI) and retrieval both propagate into concept linking. When gold entities are provided, accuracy improves substantially (e.g., GPT-4o from 0.1052 to 0.1829; DeepSeek-V3 from 0.1074 to 0.1889), confirming that entity identification is a key bottleneck in this context. With gold candidates guaranteed in the pool (rerank-only), performance further improves (e.g., GPT-4o 0.2023; DeepSeek-V3 0.2144), showing that narrowing the label space alleviates but does not eliminate difficulty, as semantic ambiguity among fine-grained US-GAAP concepts remains challenging.
Table 9. Accuracy comparison across the full pipeline, FinCL subtask, and rerank-only settings.
Model Pipeline FinCL Rerank-only
GPT-4o 0.1052 0.1829 0.2023
Deepseek-V3 0.1074 0.1889 0.2144
Llama-4-Scout-17B-16E-Instruct 0.1045 0.1649 0.1792
Llama-3.3-70B-Instruct 0.0665 0.1318 0.1482
DeepSeek-R1-Distill-Qwen-32B 0.0814 0.1141 0.1258
Qwen3-32B 0.1127 0.1277 0.1402
gemma-2-27b-it 0.0533 0.1099 0.1210
Qwen3-14B 0.1069 0.1144 0.1321
Llama-3.1-8B-Instruct 0.0575 0.0913 0.1076
Llama-3.2-3B-Instruct 0.0178 0.0415 0.0566
Qwen3-1.7B 0.0598 0.0735 0.0880
Qwen3-0.6B 0.0362 0.0414 0.0533
Fino1-8B 0.0419 0.0807 0.0908
4.5. Error cases
To better illustrate the semantic ambiguity challenge, we present a representative error case from our best-performing model, DeepSeek-V3. Consider the following input paragraph:
“Cash equivalents include term deposits with banks, money market funds, and all highly liquid investments with original maturities of three months or less…… At December 28, 2024, we had restricted cash of \$31 million recorded in other current assets and restricted cash of \$121 million recorded in other non-current assets.”
This paragraph contains 2 “monetary entities”: 31 and 121. The gold concepts and predicted concepts that are assigned to both entities are shown in Table 10. For case 1, the model’s choice was likely influenced by the phrase “recorded in current assets”. Yet the paragraph describes the overall composition of restricted cash, and the correct concept reflects this aggregate view rather than a single classification. For case 2, although both concepts denote non-current restricted cash, the model selected a narrower variant, showing the difficulty of distinguishing between highly similar concepts with subtle differences in scope.
Table 10. Error cases from DeepSeek-V3 rerank.
Case 1: \$31m
Gold Concept us-gaap:RestrictedCashAndCashEquivalentsAtCarryingValue
Gold Concept Rank 18 / 200 (not in Top-5)
Model Prediction us-gaap:RestrictedCashAndCashEquivalentsCurrent
Case 2: \$121m
Gold Concept us-gaap:RestrictedCashAndCashEquivalentsNoncurrent
Gold Concept Rank 2 / 200 (in Top-5)
Model Prediction us-gaap:CashCashEquivalentsRestrictedCashAndRestrictedCashEquivalents
5. Conclusion
@@
This paper presents 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863i0.43922 0.45098 0.95294n0.5098 0.43529 0.93725T0.58431 0.42353 0.92157a0.6549 0.40784 0.90588g0.72549 0.39216 0.8902g0.79608 0.37647 0.87451i0.87059 0.36471 0.85882n0.94118 0.34902 0.84314g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:, a benchmark for evaluating large language models on XBRL tagging of real-world financial reports. The task is divided into two subtasks, financial numeric identification (FinNI) and concept linking (FinCL), to enable fine-grained evaluation of both information extraction and taxonomy alignment. Results show that while LLMs generalize well to long-tail entities and perform competitively in zero-shot settings, they struggle with accurate alignment to GAAP concepts. This reveals limitations in structure-aware reasoning and highlights the need for better semantic understanding. 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863i0.43922 0.45098 0.95294n0.5098 0.43529 0.93725T0.58431 0.42353 0.92157a0.6549 0.40784 0.90588g0.72549 0.39216 0.8902g0.79608 0.37647 0.87451i0.87059 0.36471 0.85882n0.94118 0.34902 0.84314g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: offers a foundation for advancing research in XBRL tagging and regulatory reporting. The limitations of our work and directions for future research are discussed in Appendix A.